OpenStudy (anonymous):

Which of the following is the best linear approximation for f(x) = tan(x) near x equals 3 times pi divided by 4?

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

do you know the derivative of tangent?

OpenStudy (misty1212):

the slope will be \[\sec^2(\frac{3\pi}{4})\]

OpenStudy (misty1212):

then use the point slope formula

OpenStudy (misty1212):

the point will be \[(\frac{3\pi}{4},-1)\]

OpenStudy (anonymous):

It gives me options of 2y=4x-3pi-2 2y=2x-3pi-1 2y=4x-3pi-1 2y=x-3pi-2 @misty1212

OpenStudy (misty1212):

\[\sec^2(\frac{3\pi}{4})=2\]so you can start with \[y+1=2(x-\frac{3\pi}{4})\] and try to change to "standard form"

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

multiply out on the right then multiply all by 2

OpenStudy (misty1212):

\[y+1=2(x-\frac{3\pi}{4})\\ y+1=2x-\frac{3\pi}{2}\\ 2y+2=4x-3\pi\]

OpenStudy (misty1212):

looks like it will be choice A \[2y+4x-3\pi-2\]

OpenStudy (anonymous):

Okay I had the second line, but then I got confused from there. Thank you!

OpenStudy (misty1212):

\(\color\magenta\heartsuit\)\(\color\magenta\heartsuit\)